# Prediction Sets and Conformal Inference with

### v3

@@ Appendix C Differential conformal procedure
    Compute the relevant $(1-\alpha)(1+\frac{1}{\left\lvert\mathcal{I}_{2}\right\rvert})$ quantile $\vartheta$ of the conformity scores $\left\{s_{j}:j\in\mathcal{I}_{2}\right\}$, and the conformal prediction set is $\tilde{C}_{I}$ given by
@@
Given $\hat{C}_{S}$ estimated with the training set $\mathcal{I}_{1}$, we can construct the conformity score $s_{j}=s(Y_{j}^{L},Y_{j}^{U};\hat{C}_{S})$, where for $C\in\mathcal{C}_{M}$,
$s(y^{L},y^{U};C)=\min_{m\leq M}\max\left\{\hat{\tau}_{0m}-y^{L},y^{U}-\hat{\tau}_{1m}\right\}.$
And we can solve the following for $w=(w_{01},w_{11},w_{02},w_{12},\dots,w_{0M},w_{1M})$, $\tilde{C}_{M}=\cup_{m}[\hat{\tau}_{0m}-w_{0m},\hat{\tau}_{1m}+w_{1m}]$, by
@@ Appendix D Metrics for sets
However, when the prediction set is a union of intervals, these two metrics are not equivalent. In fact, consider $A\subset\mathbb{R}$ and $B=A\cup\left\{b\right\}$, for some $b\in A^{c}$ such that $d(b,A)>0$, then the Hausdorff distance between $A$ and $B$ is $d_{H}(A,B)=d(b,A)>0$, while the volume of the symmetric difference is 00. Thus, the volume of symmetric difference has the advantage of being less sensitive to a difference of negligible sets and seems more appropriate for practical prediction sets estimations.
@@ Appendix E Estimation and conformal inference based on quantile regression
and $\hat{C}_{q}(x)=[\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)]$ as an estimator for $C_{q}(x)=[{q^{L}_{\frac{\alpha}{2}(x)},q^{U}_{1-\frac{\alpha}{2}}(x)}]$ is an asymptotically valid, albeit not necessarily optimal, prediction set. Here $\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)$ are the consistent quantile regression estimators. There is a vast body of research on quantile regression, see Koenker, (2017), Athey et al., (2019) for random forest based quantile regression and Li et al., (2021) for nonparametric quantile regression with mixed continuous and discrete data.
Once we have $\hat{C}_{q}(x)$, it is straightforward to use the conformalisation procedure to obtain finite-sample valid prediction sets. Let $s_{j}^{q}=s(Y_{j}^{L},Y_{j}^{U},X_{j};\hat{q}_{\frac{\alpha}{2}}^{L},\hat{q}_{\frac{\alpha}{2}}^{U})$ and $\vartheta^{q}_{1-\alpha}=q(\{s_{j}^{q}:j\in\mathcal{I}_{2}\};(1-\alpha)(1+1/\left\lvert\mathcal{I}_{2}\right\rvert))$. The conformalised quantile prediction interval is

### v4

@@ Appendix C Differential conformal procedure
[…]
@@
Given $\hat{C}_{S}$ estimated with the training set $\mathcal{I}_{1}$, we can construct the conformity score $s_{j}=s(Y_{j}^{L},Y_{j}^{U};\hat{C}_{S})$, as in 4. And we can solve the following for $w=(w_{01},w_{11},w_{02},w_{12},\dots,w_{0M},w_{1M})$, $\tilde{C}_{M}=\cup_{m}[\hat{\tau}_{0m}-w_{0m},\hat{\tau}_{1m}+w_{1m}]$, by
@@ Appendix D Metrics for sets
However, when the prediction set is a union of intervals, these two metrics are not equivalent. In fact, consider $A\subset\mathbb{R}$ and $B=A\cup\left\{b\right\}$, for some $b\in A^{c}$ such that $d(b,A)>0$, then the Hausdorff distance between $A$ and $B$ is $d_{H}(A,B)=d(b,A)>0$, while the volume of the symmetric difference is 0. Thus, the volume of symmetric difference has the advantage of being less sensitive to a difference of negligible sets and seems more appropriate for practical prediction sets estimations.
@@ Appendix E Estimation and conformal inference based on quantile regression
and $\hat{C}_{q}(x)=[\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)]$ as an estimator for $C_{q}(x)=[{q^{L}_{\frac{\alpha}{2}(x)},q^{U}_{1-\frac{\alpha}{2}}(x)}]$ is an asymptotically valid, albeit not necessarily optimal, prediction set. Here $\hat{q}^{L}_{\frac{\alpha}{2}}(x),\hat{q}^{U}_{1-\frac{\alpha}{2}}(x)$ are the consistent quantile regression estimators. There is a vast body of research on quantile regression, see Koenker (2017), Athey et al. (2019) for random forest based quantile regression and Li et al. (2021) for nonparametric quantile regression with mixed continuous and discrete data.
Once we have $\hat{C}_{q}(x)$, it is straightforward to use the conformalisation procedure to obtain finite-sample valid prediction sets. Let $s_{j}^{q}=s(Y_{j}^{L},Y_{j}^{U},X_{j};\hat{q}_{\frac{\alpha}{2}}^{L},\hat{q}_{1-\frac{\alpha}{2}}^{U})$ and $)\vartheta^{q}_{1-\alpha}=q(\{s_{j}^{q}:j\in\mathcal{I}_{2}\};(1-\alpha)(1+1/\left\lvert\mathcal{I}_{2}\right\rvert$)). The conformalised quantile prediction interval is